1 and 2/5 + 3 and 2/3+ 2 and 5/6
Answer by tommyt3rd(5050)

1 2/5 + 3 2/3+ 2 5/6=
1+3+2+ 2/5 + 2/3+ 5/6=
6 + 12/30+20/30+25/30=
6 +57/30=
7 27/30=
7 9/10